In reaction

CH.(9)+30 (9) ► 20092110

How many (a)moles of O are consumed and (b)moles of H,O are formed when 4.16 102 mol of CH, react?
1
Expert's answer
2020-09-15T13:52:57-0400

From the context of this question: CH4+2O2 = CO2+2H2O

n(O2)= n (H2O) =2 x n (CH4) = 2 x 4.16102 = 8.32204 moles